OpenStudy (anonymous):

Could someone explain how to do these problems? Find the first six terms of the sequence. a1 = -3, an = 2 ● an-1

Nnesha (nnesha):

n = term that you have to find so to find 2nd term replace n by 2

OpenStudy (sohailiftikhar):

a1=-3

Nnesha (nnesha):

\[\huge\rm, a_1=3\] \[\huge\rm a_n=2 \times a_{n-1}\] substitute n for 2 \[\huge\rm a_n=2 \times a_{2-1}\] 2-1 = 1 then substitute a_1 for -3

OpenStudy (anonymous):

Ok so...would it be: -3, -6, -12, -24, -48, -96 @Nnesha

Nnesha (nnesha):

:)

OpenStudy (anonymous):

\[a_{1}=-3\]\[a_{n}=2 \times a_{n-1}\] That is to say each term is equal to twice the previous term, if you are given the first term as -3 u can calculate upto as many terms as u like, each successive term is twice the last term second term will be 2 times the first time, 3rd term will be 2 times 2nd term and so on

OpenStudy (anonymous):

looks good!

OpenStudy (anonymous):

Could you help me with another one? @Nishant_Garg

OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, 6) and a directrix at y = -6. @Nishant_Garg I never know quite where to start.

OpenStudy (anonymous):

hm I think the equation will be \[x^2=4ay\] Becuase the focus lies on the y axis, the equation must be either the form of \[x^2=4ay\] Or \[x^2=-4ay\] But since the equation of directrix has a minus sign, it implies the the equation should be the first one  The first figure makes sense because the directrix is suppose to be like in front the parabola Also you are given \[a=6\]

OpenStudy (anonymous):

so just plug that a into the equation

OpenStudy (anonymous):

Thank you!